two metallic charge spheres wbhose radii are 20cm and 10cm respectively, have each 150mC charge. the common potential after they are connected by a conducting wire is


1
Expert's answer
2021-07-08T13:23:56-0400

Given quantities:

"R_1 = 20cm = 0.2m"

"R_2 = 10cm = 0.1m"

"q_1=q_2=150mC=15*10^{-2}C"

"Common\space potential = \large\frac{Total \space Charge}{Total\space capacitance}=\large\frac{q_1+q_2}{C_1+C_2}=\frac{q_1+q_2}{4\pi e_o(R_1+R_2)} = \frac{150*10^{-6}*2}{4*3.14*8.85*10^{-12}*0.3}" = "9*10^6V"